OpenStudy (anonymous):

An object travels through a medium so that its position s after t seconds is s(t) =10 - (5/radical(t)). To the nearest thousandth, evaluate the average velocity of the object on the intervals [ 4, 5 ],[ 4, 4.5 ], [ 4, 4.1 ], [ 4, 4.05 ]. How do I do atleast one interval?

OpenStudy (anonymous):

find the slope

OpenStudy (anonymous):

for example, the first one is \[\frac{s(5)-s(4)}{5-4}\] all the others are the same, you need a calculator

OpenStudy (anonymous):

\[\frac{s(4.5)-s(4)}{4.5-4}\] \[\frac{s(4.1)-s(4)}{4.1-4}\] \[\frac{s(4.05)-s(4)}{4.05-4}\] a spread sheet would make this very easy

OpenStudy (anonymous):

but it is a numerical exercise. first one for example is \[\frac{1-\frac{5}{\sqrt{5}}}{1}=1-\sqrt{5}=-1.236\]

OpenStudy (anonymous):

That's what I did and I got the first one wrong.. I must have hit something wrong... I'll try again. Thanks!